Consider two assets: Unicorn stock and Mermaid stock. Unicorn stock either provides a rate of return of 30% or -20%, with equal probability. Mermaid stock either provides a return of 15% or -5%, with equal probability. You decide to hold a portfolio by investing half of your money in Unicorn stock and the other half in Mermaid stock. If Unicorn stock and Mermaid stock’s returns are independent of each other, what will be the expected return and standard deviation of your portfolio? Assume the risk-free interest rate is 3%. Do you think you have made a rational choice by holding the portfolio? 


1
Expert's answer
2021-02-04T10:06:36-0500

My expected return is ((30+(-20)/2)×0.5+((15+(-5)/2)×0.5=5. Standart deviation is (30-5)×0.5+(15-5)×0.5=17.5.Risk-free interest rate is 3%, but holding the portfolio l can get 5-17.5=-12.5%, so it is unreasonable